Find the smallest integral value of c for which the roots of 2x^2-8x+C=0 ARE imaginary.... please help and thanks
Answer by stanbon(75887)

Find the smallest integral value of c for which the roots of 2x^2-8x+C=0 ARE imaginary....
---------------
The roots are imaginary when the discriminant is negative.
a = 2 ; b = -8 ; c = C
------
discriminant = b^2-4ac = (-8)^2 - 4*2*C < 0
64 - 8C < 0
8C > 64
C > 8
----------------
The roots are imaginary when C > 8.
=====================================
Cheers,
Stan H.